S=-120 + 0.25y investment level is $400 determine the equilibrium level using the savings/investment approach to equilibrium.


1
Expert's answer
2021-12-02T10:40:50-0500

Solution:

Under this approach, equilibrium is where: S = I

-120 + 0.25Y = 400

0.25Y = 400 + 120 = 520

0.25Y = 520

Y = 2,080


The equilibrium level = 2,080.